A construction company signed a loan contract at 4.18​% compounded annually​, with the provision to pay ​$455 at the end of each month for three years.

​(a) What is amount of the​ loan?

​(b) How much will be owed at the end of twenty ​months?

​(c) How much of the principal will be repaid within the first twenty ​months?

​(d) How much interest is paid during the first twenty months?
Answer #1

The formula is,

a. Monthly payment=Loan amount*monthly interest rate*((1+monthly interest rate)^n)/[((1+monthly interest rate)^n)-1]

monthly payment=455

monthly interest rate=4.18%/12=0.3483%

n=total number of months=3*12=36

455=loan amount*0.3483%*((1+0.3483%)^36)/[((1+0.3483%)^36)-1]

loan amount=(455/0.3483%)*[((1+0.3483%)^36)-1]/((1+0.3483%)^36)

loan amount=130622*0.133355/1.133355=$153,69.48

Please find the below amortization schedule with formulas, how I arrived at

b. The outstanding balance at the end of twenty months=$7068.80

c. The total principal paid in twenty months=$8300.68

d. The total interest paid in twenty months=$799.32

Periods Opening Balance Monthly payment Interest amount=(Opening Balance*0.3483%) Principal amount=monthly payment-Interest Ending Balance=Opening Balance-Principal
1 15369.48 455.00 53.53 401.47 14968.01
2 14968.01 455.00 52.13 402.87 14565.15
3 14565.15 455.00 50.73 404.27 14160.88
4 14160.88 455.00 49.32 405.68 13755.20
5 13755.20 455.00 47.91 407.09 13348.11
6 13348.11 455.00 46.49 408.51 12939.60
7 12939.60 455.00 45.07 409.93 12529.67
8 12529.67 455.00 43.64 411.36 12118.31
9 12118.31 455.00 42.21 412.79 11705.52
10 11705.52 455.00 40.77 414.23 11291.29
11 11291.29 455.00 39.33 415.67 10875.61
12 10875.61 455.00 37.88 417.12 10458.49
13 10458.49 455.00 36.43 418.57 10039.92
14 10039.92 455.00 34.97 420.03 9619.89
15 9619.89 455.00 33.51 421.49 9198.40
16 9198.40 455.00 32.04 422.96 8775.43
17 8775.43 455.00 30.56 424.44 8351.00
18 8351.00 455.00 29.09 425.91 7925.09
19 7925.09 455.00 27.60 427.40 7497.69
20 7497.69 455.00 26.11 428.89 7068.80
799.32 8300.68